# ZooplanktonBench: A Geo-Aware Zooplankton Recognition and Classification Dataset from Marine Observations

### v1

@@ Abstract.
Plankton are small drifting organisms found throughout the world’s oceans. One component of this plankton community is the zooplankton, which includes gelatinous animals and crustaceans (e.g. shrimp), as well as the early life stages (i.e., eggs and larvae) of many commercially important fishes. Being able to monitor zooplankton abundances accurately and understand how populations change in relation to ocean conditions is invaluable to marine science research, with important implications for future marine seafood productivity. While new imaging technologies generate massive amounts of video data of zooplankton, analyzing them using general-purpose computer vision tools turns out to be highly challenging due to the high similarity in appearance between the zooplankton and its background (e.g., marine snow). In this work, we present the ZooplanktonBench, a benchmark dataset containing images and videos of zooplankton associated with rich geospatial metadata (e.g., geographic coordinates, depth, etc.) in various water ecosystems. ZooplanktonBench defines a collection of tasks to detect, classify, and track zooplankton in challenging settings, including highly cluttered environments, living vs non-living classification, objects with similar shapes, and relatively small objects. Our dataset presents unique challenges and opportunities for state-of-the-art computer vision systems to evolve and improve visual understanding in a dynamic environment with huge variations and be geo-aware. The code and settings described in this paper can be found on our website: https://lfk118.github.io/ZooplanktonBench_Webpage.
@@ 4.1.1. Fine-grained detection and classification on zooplankton species
One of the unique challenges in our data/task is the predominance existence of distractors (i.e., non-living marine snow particles) in the data. As shown in Figure 4, there are a lot of objects that are visually similar to the zooplankton but they are actually belonging to “marine snow.” The high similarity in their appearances highlights the difficulty of differentiating living zooplankton from marine snow. The marine snow consists of biological debris and organic material falling from upper waters to the deep ocean (NOAA, [n. d.]). To make matters more challenging, in the video data, not only zooplankton can move in the water, the marine snow objects can also move in motion due to the disturbance of the water around them.
@@ 5.3. Large visual-language model
We also design a second prompt to explicitly encode geo-metadata to evaluate the performance of GPT-4V being geo-aware: “You are a Marine Biologist. I am sending you an image containing zooplankton. The image was captured at latitude 29.5063167 and longitude 88.5215167, situated 34.5 meters beneath the seafloor. The environmental conditions at this location were as follows: an irradiance of 1.70E-05, a water pressure of 34.73 dbar, a temperature of 22.21 degrees Celsius, a salinity of 36.36 psu, and a dissolved oxygen level of 1.8678 mg per L. Given these conditions, the possible classes of zooplankton are: chaetognath, larval fish, hydromedusa, lobate ctenophore, pleurobrachia, shrimp, siphonophore, stomatopod larva, unknown, thaliac, polychaete worm, scyphomedusa, other ctenophore. Please give me the most probable name shown by the given image in the following format: class name. Please just give a short answer and do not provide explanations for your choice.“

### v3

@@ Abstract.
Plankton are small drifting organisms found throughout the world’s oceans and can be indicators of ocean health. One component of this plankton community is the zooplankton, which includes gelatinous animals and crustaceans (e.g. shrimp), as well as the early life stages (i.e., eggs and larvae) of many commercially important fishes. Being able to monitor zooplankton abundances accurately and understand how populations change in relation to ocean conditions is invaluable to marine science research, with important implications for future marine seafood productivity. While new imaging technologies generate massive amounts of video data of zooplankton, analyzing them using general-purpose computer vision tools turns out to be highly challenging due to the high similarity in appearance between the zooplankton and its background (e.g., marine snow). In this work, we present the ZooplanktonBench, a benchmark dataset containing images and videos of zooplankton associated with rich geospatial metadata (e.g., geographic coordinates, depth, etc.) in various water ecosystems. ZooplanktonBench defines a collection of tasks to detect, classify, and track zooplankton in challenging settings, including highly cluttered environments, living vs non-living classification, objects with similar shapes, and relatively small objects.Our dataset presents unique challenges and opportunities for state-of-the-art computer vision systems to evolve and improve visual understanding in dynamic environments characterized by significant variation and the need for geo-awareness. The code and settings described in this paper can be found on our website: https://lfk118.github.io/ZooplanktonBench_Webpage.
@@ 4.1.1. Fine-grained detection and classification on zooplankton species
One of the unique challenges in our data/task is the predominance existence of distractors (i.e., non-living marine snow particles) in the data. As shown in Figure 4, there are a lot of objects that are visually similar to the zooplankton but they are actually belonging to “marine snow.” The high similarity in their appearances highlights the difficulty of differentiating living zooplankton from marine snow. The marine snow consists of biological debris and organic material falling from upper waters to the deep ocean (Oceanic and Administration, 2020). To make matters more challenging, in the video data, not only zooplankton can move in the water, the marine snow objects can also move in motion due to the disturbance of the water around them.
@@ 5.3. Large visual-language model
We also design a second prompt to explicitly encode geo-metadata to evaluate the performance of GPT-4V being geo-aware (Mac Aodha et al., 2019; Mai et al., 2020; Yang et al., 2022; Mai et al., 2023a, b; Cole et al., 2023; Wu et al., 2024): “You are a Marine Biologist. I am sending you an image containing zooplankton. The image was captured at latitude 29.5063167 and longitude 88.5215167, situated 34.5 meters beneath the seafloor. The environmental conditions at this location were as follows: an irradiance of 1.70E-05, a water pressure of 34.73 dbar, a temperature of 22.21 degrees Celsius, a salinity of 36.36 psu, and a dissolved oxygen level of 1.8678 mg per L. Given these conditions, the possible classes of zooplankton are: chaetognath, larval fish, hydromedusa, lobate ctenophore, pleurobrachia, shrimp, siphonophore, stomatopod larva, unknown, thaliac, polychaete worm, scyphomedusa, other ctenophore. Please give me the most probable name shown by the given image in the following format: class name. Please just give a short answer and do not provide explanations for your choice.“